OpenStudy (gabylovesyou):

Let f(x) = twenty-four over the quantity of 3x minus two. Find f(-2).

OpenStudy (gabylovesyou):

@blurbendy

OpenStudy (gabylovesyou):

@amistre64

OpenStudy (blurbendy):

f(x) = 24 / (3x -2) f(-2) = 24 / (3(-2) - 2) f(-2) = 24 / (-6 - 2) = 24 / -8 f(-2) = -3

OpenStudy (blurbendy):

-3

OpenStudy (gabylovesyou):

thank you! ... i thought i saw -24 cx

OpenStudy (blurbendy):

ah, yup. gotta be careful w/ signs